OpenStudy (anonymous):

I need help with an AP Statistics simulation, here goes: Consider the following problem. An airline has 40 seats, and the airline accepts 43 reservations for a certain flight. If the airline's records indicate a no-show rate for this flight of 10 percent, conduct a simulation to find the following: (a) The expected number of empty seats and (b) the expected number of ticketed passengers who do not get a seat

OpenStudy (amistre64):

43*10% = number of expected no shows id think

OpenStudy (aravindg):

new question

OpenStudy (aravindg):

plz hlp

OpenStudy (amistre64):

and 3/40 should be the expected number of passangers who by a ticket and dont get a seat

OpenStudy (amistre64):

well ... i think that amounts to 3 passengers expected not to get a seat lol

OpenStudy (amistre64):

43*.1 = 4.3 noshows; out of 43 tickets; say 5 43 - 5 = 38 40 - 38 = 2 empty seats